Two angles are complimentary if they add up to 90* if one angle is (4x+5)* and the other angle is (2x+7)* find x
Found 2 solutions by harpazo, stanbon:
Answer by harpazo(655)

4x+5) + (2x+7) = 90
Solve for x.

Answer by stanbon(75887)

Two angles are complimentary if they add up to 90* if one angle is (4x+5)* and the other angle is (2x+7)* find x
---------
Equation:
4x+5 + 2x+7 = 90
-----
6x + 12 = 90
6x = 78
x = 13
===================
4x+5 = 57 degrees
2x+7 = 33 degrees
=====================
Cheers,
Stan H.
=====================